OpenStudy (lerenstudy):

In this exercise, find the constant a such that the function is continuous on the entire real line.

OpenStudy (brainzonly):

.....

OpenStudy (brainzonly):

I don't get it. :P Is there more?

OpenStudy (lerenstudy):

\[f(x)= 3x^2 when x \ge1 , ax-4 when x <1\]

OpenStudy (brainzonly):

ok. just a minute.

OpenStudy (lerenstudy):

Sorry, I couldn't get it to look very good, and I figured it might be easier to understand with at least the proper signs. I solved it and I got a=-1 but I read somewhere that they got a=7 which is why I was confused.

OpenStudy (brainzonly):

The only point that could be a problem is at x=3 where the step is. At x=3, f(x)=x^3=27. The limit of f as x approaches 3 from the right needs to be the same. 3^2=9. You need a*9=27, a=3

OpenStudy (brainzonly):

fan me by clicking on my pic so you can send me a PM if you have any more questions. I've got a ton of people tagging me :D Bye

OpenStudy (lerenstudy):

Okay, so I kind of get that, but why wouldn't you use 1, since that is mentioned in the piecewise?

OpenStudy (lerenstudy):

Because the way I solved this was looking at whether or not the lim 3x^2 as x approaches 1 from the left is the same as lim ax-4 as x approaches 1 from the right

OpenStudy (brainzonly):

hmmm...

OpenStudy (lerenstudy):

Since in order for a function to be continuous both functions must be equivalent (meaning that they meet at the same point)

OpenStudy (brainzonly):

you're right

OpenStudy (lerenstudy):

And from there, substitution could be done to find the a value so that it equals the same as what the limit of 3x^2 is

OpenStudy (brainzonly):

I think you lost me on that one :P

OpenStudy (brainzonly):

wait I get it!

OpenStudy (brainzonly):

I think you were right the 1st time...

OpenStudy (lerenstudy):

So then the answer would be a=-1 because this causes the two limits to be equivalent?

OpenStudy (lerenstudy):

There's a little problem in that the back of the book says that the answer is indeed 7, which doesn't quite make sense for me. However, another problem I did, I got the answer right by the same process.

OpenStudy (lerenstudy):

I see what I did wrong. I thought it was +4 instead of -4